OpenStudy (anonymous):

Simplify √3y divided by √y a)1/3 b)√3 c)3

OpenStudy (anonymous):

The product of the square roots is the same as the square root so: \[\sqrt{3y}=\sqrt{3}\sqrt{y}\] The sqrt(y) will cancel so the answer is \[\sqrt{3}\]